# UnlearnCanvas: A Stylized Image Dataset to Benchmark Machine Unlearning for Diffusion Models

### v2

@@ 6 Experiment Results
Understanding unlearning method’s behavior via unlearning directions. As noted earlier, different unlearning methods display distinct unlearning behaviors. To gain insights into the underlying reasons for these differences, Fig. 8 (a) and (b) visualize the ‘unlearning directions’ for ESD and SalUn, respectively. These unlearning directions are determined by connecting the unlearning target with the predicted label of the generated image from the unlearned DM conditioned on the unlearning target. As shown in Fig. 8 (a), ESD demonstrates a focused shift in image generation after object unlearning, with a predominant transition towards generating images labeled by ‘Sea’ and ‘Trees’. This behavior arises from ESD’s optimization process, designed to steer the generation of the DM away from a predefined concept. Consequently, images generated by the ESD-induced unlearned model consistently lack clearly identifiable objects, resembling waves and trees, which leads to their classification into the ‘Sea’ and ‘Trees’ classes; see Fig. 8 (c) for examples of generated images. In contrast, SalUn exhibits a more diverse range of unlearning directions, shifting images to 11 different objects. This diversity results from SalUn’s requirement to replace the unlearning target with a random concept. As shown in Fig. 8 (c), images generated by SalUn post-object unlearning still maintain some object contours (different from the original unlearning target) and better retain style information compared to ESD.
Expanding applications to style transfer. In addition to evaluating DMs’ unlearning performance, it’s worth noting that UnlearnCanvas has the potential to extend its utility to another generative modeling task: style transfer (Liu et al., 2021; Huang & Belongie, 2017; Li et al., 2019; Chen & Schmidt, 2016; Park & Lee, 2019; An et al., 2021). One significant challenge in evaluating style transfer methods is the absence of ground truth images for the given reference style image and content image (Cai et al., 2023). Thus, quantitative evaluations are also very limited in the literature. In Appx. C, we utilize UnlearnCanvas to assess 9 latest style transfer methods, providing enriched and precise evaluations. While not the primary focus of this work, this illustrates the great potential of UnlearnCanvas to contribute broader visual generative modeling tasks.
7 Conclusion
In this paper, we systematically reviewed the existing the stateful MU (machine unlearning) methods on DMs (diffusion models), and pinpointed the key challenges rooted in their evaluation systems, which could lead to incomplete, inaccurate, and biased assessments. In response, we propose UnlearnCanvas, a high-resolution stylized image dataset facilitating to build a comprehensive evaluation system for MU methods, together with a novel systematic retainability measurements, that have been long overlooked but are essential in a all-round evaluation. By benchmarking 5 stateful MU methods, we reveal novel insights into the their strengths and weaknesses, deepen the understanding on the underlying mechanism, and further illuminate several promising research directions. In the end, with a case study on style transfer, we show UnlearnCanvas enjoys high versatility and great potential to benchmark other generative modeling tasks beyond MU.
Impact Statement
$\bullet$ Avoiding Copyright Issues. There is an urgent need for the generative model providers to scrub the influence of certain data on an already-trained model. In January 2023, a notable lawsuit targeted two leading AI art generators, Stable Diffusion (Rombach et al., 2022) and Midjourney (Midjourney et al.,, 2023), for alleged copyright infringement. Concurrently, incidents with the recently released Midjourney V6 (Midjourney et al.,, 2023) also highlighted a visual plagiarism issue on famous film scenes. These instances illuminate the broad copyright challenges inherent in the way of training data collection method of those foundation generative models’ training datasets. MU methods can be used as an effective method to remove the influence of the private data and avoid unnecessary retraining.
$\bullet$ Mitigating biases and stereotypes. Generative AI systems are known to have tendencies towards bias, stereotypes, and reductionism, when it comes to gender, race and national identities (Turk, 2023). For example, a recent study on the images generated with Midjourney revealed, that images associated with higher-paying job titles featured people with lighter skin tones, and that results for most professional roles were male-dominated (Vinker et al., 2023). MU is known to be effective in eliminating biases rooted in the training data. Moreover, UnlearnCanvas offers a flexible framework to benchmark MU techniques against bias removal, allowing for the creation and quantitative control of biases across different object classes for comprehensive bias removal studies.
@@ 7 Conclusion
A Detailed Experiment Settings
A.1 Fine-tuning Style and Object Classifier with UnlearnCanvas.
Style and object classifiers need to be trained as part of the test bed proposed in our evaluation pipeline (Fig. 5). Here, we adopted a pretrained ViT-L/16 model (Dosovitskiy et al., 2020) pretrained on ImageNet and fine-tune it on UnlearnCanvas. UnlearnCanvas are split into the train set and test set with a ratio of $9:1$. After hyper-parameter tuning, the classifiers are trained with Adam optimizer with a learning rate of $0.01$ for 10 epochs. We noticed the model training accuracy can readily converge to $100\%$ even with 10 epochs of training.
A.2 Fine-tuning Stable Diffusion Model with UnlearnCanvas.
@@ A Detailed Experiment Settings
Practically, we fine-tune the pretrained StableDiffusion v1.5 on UnlearnCanvas for 20k steps with a learning rate of $61e-$ 6. Unless otherwise stated, we strictly follow the training configurations used in Stable Diffusion (Rombach et al., 2022). For each image in UnlearnCanvas, we annotate the data with text prompt An image of {object} in {style} style, where the object and style are the corresponding object and style label. For seed images, the style label we use is ‘photo’ style. We use the training scripts provided by Diffuser official tutorial (StableDiffusion) and the pretraining model card is runwayml/stable-diffusion-v1-5. During training, the checkpoints will be saved every 1000 steps.
@@
To evaluate the quality of the saved checkpoints and selcect the best one for unlearning study, the checkpoints are first used to generate an image set with the same prompt as training (An image of {object} in {style} style) by traversing all the possible style and object labels. Each prompt are used to generate $5$ images with different random seed. The image set for each checkpoint are fed into the style and object classifier trained in Appx. A.1. The model with the highest average performance on all the styles and objects are selected as the testbed for MU study. The classification performance are also used as a reference for later IRA/CRA comparison, which are disclosed in the first row of Fig. 7 (left). In Sec. 6, the experiments were built on the $50$ out of the $60$ styles, which served as the first edition of the dataset, and these styles are listed in Fig. 7. We remark that similar experiments can be done on all the 60 styles.
A.3 Implementation of the Tested Machine Unlearning Methods
In this work, we inspected 5 stateful MU methods for DMs. For each method, we use their publicly released source codes as code bases, which are listed below:
@@ A.3 Implementation of the Tested Machine Unlearning Methods
  • ESD: based on the suggestions from the authors, ESD is used to only fine-tune the cross attention-related model weights (ESD-x). Other settings strict follow the ones used in the paper.
@@
  • UCE: This method requires a guided concept (prompt) for each unlearning concept. For style unlearning, we use the prompt An image in {style*} style as the guided concept, where style* represents the next style in UnlearnCanvas in alphabetical order. Similarly, An image of object* is used for object unlearning, where object* is the next object in UnlearnCanvas in alphabetical order.
@@
B Additional Machine Unlearning Experiment Results

### v4

@@ 4.1 Benchmarking Current DM Unlearning Methods for Style and Object Unlearning
As noted earlier, different unlearning methods display distinct unlearning behaviors. To gain insights into the underlying reasons for these differences, Fig. 7 (a) and (b) visualize the ‘unlearning directions’ for ESD and SalUn, respectively. These unlearning directions are determined by connecting the unlearning target with the predicted label of the generated image from the unlearned DM conditioned on the unlearning target. As shown in Fig. 7 (a), ESD demonstrates a focused shift in image generation after object unlearning, with a predominant transition towards generating images labeled by ‘Sea’ and ‘Trees’. This behavior arises from ESD’s optimization process, designed to steer the generation of the DM away from a predefined concept. Consequently, images generated by the ESD-induced unlearned model consistently lack clearly identifiable objects, resembling waves and trees, which leads to their classification into the ‘Sea’ and ‘Trees’ classes; see Fig. 7 (c) for examples of generated images. In contrast, SalUn exhibits a more diverse range of unlearning directions, shifting images to 11 different objects. This diversity results from SalUn’s requirement to replace the unlearning target with a random concept. As shown in Fig. 7 (c), images generated by SalUn post-object unlearning still maintain some object contours (different from the original unlearning target) and better retain style information compared to ESD.
Figure 8: UA of DM unlearning against adversarial prompts [59]. Unlearned models in Tab. 2 are used as victim models to generate adversarial prompts.
4.2 Benchmarking Current DM Unlearning Methods in More Challenging Scenarios
Unlearning robustness against adversarial prompts.
Recent studies [59, 60] have highlighted the vulnerabilities of DM unlearning to adversarial prompts, such as jailbreak attacks. In Fig. 8, we use the state-of-the-art attacking method, UnlearnDiffAtk [59], to craft adversarial prompts based on UnlearnCanvas and evaluate the robustness of unlearned DMs shown in Tab. 2. See Appx. B.4 for detailed setup. As we can see, all the DM unlearning methods experience a significant drop in UA, falling below $60\%$ when confronted with adversarial prompts. Notably, methods like UCE, SEOT, and SHS exhibit particularly steep declines. Moreover, a high UA under normal conditions does not necessarily imply robustness to adversarial attacks. For instance, although UCE achieves a UA over 90% in normal settings, its performance plummets to below 50% against adversarial prompts. This stark contrast highlights the importance of the worst-case evaluation for MU methods.
Table 3: Performance of unlearning style-object combinations. The assessment includes UA and retainability in three contexts: SC (style consistency), OC (object consistency), and UP (unrelated prompting). The best result in each metric is highlighted in green.
Method UA SC OC UP
ESD [23] $91.42\%$ ${4.88\%}$ ${14.72\%}$ $84.38\%$
FMN [28] ${45.37\%}$ $68.73\%$ $62.74\%$ $83.25\%$
UCE [24] $75.97\%$ ${4.53\%}$ ${5.72\%}$ ${35.42\%}$
CA [25] ${47.92\%}$ ${10.08\%}$ ${56.35\%}$ $81.54\%$
SalUn [27] ${42.21\%}$ $62.45\%$ $70.93\%$ $87.28\%$
SEOT [30] ${29.32\%}$ ${45.31\%}$ ${53.64\%}$ $85.45\%$
SPM [26] ${45.72\%}$ ${41.34\%}$ ${36.32\%}$ $67.82\%$
EDiff [31] $71.33\%$ ${35.23\%}$ ${26.32\%}$ ${51.52\%}$
SHS [32] ${55.32\%}$ ${14.34\%}$ ${24.32\%}$ $83.95\%$
Unlearning at a finer scale: Style-object combinations as unlearning targets.
Next, we consider the fine granularity of the unlearning target, defined by a style-object combination, such as “An image of dogs in Van Gogh style”. This unlearning challenge requires the unlearned model to avoid affecting image generation for dogs in non-Van Gogh styles and Van Gogh-style images with non-dog objects. See Appx. B.5 for detailed setup. Here, besides UA, the retainability performance is assessed in three contexts. (a) Style consistency (SC): retainability under prompts with different objects but in the same style, e.g., “An image of cats in Van Gogh style”. (b) Object consistency (OC): retainability under prompts featuring the same object in different styles, e.g., “An image of dogs in Picasso style”. (c) Unrelated prompting (UP): retainability for all other prompts. Tab. 2 presents the performance of unlearning style-object combinations. Style-object combinations are more challenging to unlearn than individual objects or styles, as evidenced by a significant drop in UA—over 20% lower compared to values in Tab. 2. Retainability drops to below 20% for top-performing methods like ESD and UCE, originally highlighted for their efficacy. This is presumably due to ESD’s underlying unlearning mechanism, which requires only a single prompt, resulting in a poor ability to precisely define the unlearning scope.
Evaluation in sequential unlearning.
Furthermore, we evaluate the unlearning performance in sequential unlearning (SU) [70, 71], where unlearning requests arrive sequentially. This parallels the continual learning (CL) task, which requires models to not only unlearn new targets effectively but also maintain the unlearning of previous targets while retaining all other knowledge. Here, we consider unlearning 6 styles sequentially and the results are presented in Tab. A5. We remark that the method SEOT does not support sequential unlearning in its original implementation and thus is not included in Tab. A5. Our findings reveal significant insights. (1) Degraded retainability: Sequential unlearning requests generally degrade retainability across all methods, with RA values frequently dropping below the average levels previously seen in Tab. 2. Here RA is given by the average of IRA and CRA. (2) Unlearning rebound effect: Knowledge previously unlearned can be inadvertently reactivated by new unlearning requests. This is evidenced by decreasing UA values for earlier objectives as more unlearning tasks are introduced. This suggests that residual knowledge remains within the model and can be reactivated, aligning with findings from Fig. 8. This indicates the unlearned models by some MU methods do not essentially lose the generation ability of the unlearning target. (3) Catastrophic retaining failure: RA significantly drops at a certain request, exemplified by a sudden decrease in RA of UCE from $81.42\%$ to $29.38\%$ after the second request, $\mathcal{T}_{2}$. This indicates that the seemingly acceptable side effects generated by some unlearning methods will drastically modify the knowledge representations when accumulated. This experiment illuminates the complex dynamics of knowledge removal and retention within DMs and highlights the potential pitfalls of existing unlearning methods when faced with sequential unlearning tasks. The observation of the ‘unlearning rebound effect’ and ‘catastrophic retaining failure’ particularly emphasizes the need for a more nuanced understanding of how knowledge is managed within DMs.
Visualizations.
We provide plenty of visualizations, illustrations, and qualitative results. These visualizations are intended to deepen the understanding of the effects of different MU methods and clearly illustrate the challenges identified in previous sections. Specifically, in Fig. A11, we provide abundant generation examples of all the 9 methods benchmarked in this work in a case study of unlearning the ‘Cartoon’ style. Both the successful and failure cases are demonstrated in the context of unlearning effectiveness, in-domain retainability, and cross-domain retainability. Besides, in Fig. A12, we provide visualizations for the effect of adversarial prompts enabled by UnlearnDiffAtk [59].
5 Conclusion, Discussion, and Limitation
In this paper, we systematically reviewed existing MU (machine unlearning) methods on DMs (diffusion models) and identified key challenges in their evaluation systems that could lead to incomplete, inaccurate, and biased assessments. In response, we propose UnlearnCanvas, a high-resolution stylized image dataset designed to facilitate comprehensive evaluation of MU methods. We also introduce novel systematic evaluation metrics. By benchmarking nine state-of-the-art MU methods, we reveal novel insights into their strengths and weaknesses and also deepen the understanding of their underlying mechanisms. Our analysis of three challenging unlearning tasks highlights significant shortcomings of current methods, including a lack of robustness, difficulties in fine-scale unlearning, and issues arising from sequential unlearning tasks. These findings also point to meaningful future research directions aimed at developing satisfactory and practical MU methods for real-world applications. Furthermore, we recognize limitations in our benchmark, such as its focus on specific Stable Diffusion models and the text-to-image task in DMs. We hope this work serves as a foundation for future research to broaden MU evaluations across a wider range of model architectures and tasks in DMs.
Acknowledgement
The work of Y. Zhang, C. Fan, Y. Zhang, Y. Yao, J. Jia, J. Liu, and S. Liu was supported by the Cisco Research Award, the ARO Award W911NF2310343, the National Science Foundation (NSF) CISE Core Program Award IIS-2207052, and the NSF CAREER Award IIS-2338068, and the Amazon Research Award for AI in Information Security.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] The introduction (Sec. 1) clearly specifies the scope of this paper and the major contributions.
    2. (b)
      Did you describe the limitations of your work? [Yes] The limitations are discussed in Sec. 5
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [No] This work aims to improve the evaluation of machine unlearning technique, which will bring many positive societal impacts, such as mitigating copyright infringement, avoiding sexual or violent content generation, and mitigating the biases and stereotypes of the diffusion models. Given the technical focus of this work, there are no specific negative societal consequences directly stemming from it that need to be highlighted here.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A] There are no theoretical results in this paper.
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A] There are no theoretical results in this paper.
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] There are detailed instructions of the codes in the github repo to facilitate reproducing the results. The detailed experiment settings are also disclosed in Appx. B.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] Detailed training settings are disclosed in Sec. 4 and Appx. D.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No] The results reported in this work were averaged over dozens of individual unlearning cases, and the variance confirmed to be below 1%. We believe this indicates negligible variability. Given the large scale of the experiments, error bars were omitted to enhance readability and maintain focus on the key findings.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] The compute resources are disclosed in Appx. B.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] The seed images used to curate UnlearnCanvas belong to the existing assets, and their sources are properly cited in Sec. 3.
    2. (b)
      Did you mention the license of the assets? [Yes]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] This paper includes a new dataset UnlearnCanvas and the link is released both in the abstract as well as in Appx. A.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] The data being used is collected from an open-source photography website and this is mentioned in Sec. 3.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] The data does not contain any personally identifiable information or offensive content, as discussed in Appx. A.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A] No human subjects are involved.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A] No human subjects are involved.
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A] No human subjects are involved.
@@ Acknowledgement
[…]
1
A UnlearnCanvas Dataset Details
In this section, we provide a detailed description of the dataset following ‘datasheet for datasets’ [72].
A.1 Motivation
This dataset is collected to enable a precise, comprehensive, and automated quantitative evaluation framework for MU (machine unlearning) methods in DMs (diffusion models). The current evaluation plans used in the existing literature have exposed several weaknesses, which may lead to incomplete, inaccurate or even biased results, see a more detailed discussion in Sec. 2. To the best of our knowledge, there are no datasets specifically designed to meet the assessing requirements of DM unlearning. Therefore, UnlearnCanvas is designed, collected, and made to fill in this gap.
A.2 Composition: Styles and Object Classes
There are 60 predetermined artistic styles provided by Fotor [11]. The images in the same artistic style all share high stylistic consistency, which enables a high-precision style classifier to be trained on them. In Fig. A9, we list some examples of the images in each style to illustrate these styles. There are 20 distinct object classes in UnlearnCanvas. In Fig. A10, we list some examples of the images in each object to illustrate these classes.
A.3 Collection and Labeling Process
The construction of UnlearnCanvas involves two main steps: seed image collection and subsequent image stylization; see Fig. 3 for an illustration. For seed image collection, a set of high-resolution real-world photos are collected from Pexels [69], providing open-sourced photographs. There are 20 seed images collected for each of the 20 object classes; see Fig. A10. After collecting the seed images, we stylize each and every seed image into all 60 predetermined artistic styles; see Fig. A9 with Fotor [11]. After the stylization of all the images, the dataset is structured in a hierarchical manner, and each image is labeled with both its style and object classes. In order to support text-to-image training, each image is annotated with the prompt ‘An image of object in style style.’.
A.4 Uses
UnlearnCanvas can be used to evaluate MU methods in different unlearning scenarios. Please see Sec. 4 for more details. In addition, we stress that UnlearnCanvas can be used for more real-world tasks than unlearning, and we provide an example of how UnlearnCanvas can be used to systematically evaluate another generative task, style transfer, in Appx. F. We provide very detailed instructions with codes in the GitHub code repository.
A.5 Distribution
UnlearnCanvas is an open-sourced dataset and is based on the existing open-sourced data [69]. We make access to the dataset public under the MIT license. We remark that no personally identifiable information or offensive content is included in this dataset. The dataset can be accessed either through Google Drive and HuggingFace. More resources on the dataset, such as the introduction video and the benchmark, can be found in the official project webpage.
A.6 Maintenance
The dataset will be maintained by the lead author Yihua Zhang. If needed, the email for contacting is zhan1908@msu.edu. The dataset may be updated if needed (with the inclusion of more seed images, more artistic styles, and more objects). The updates will be ad-hoc and will not be periodical. Each time the dataset is updated, the updates will be reflected in the same GitHub code repository.
A.7 Author Statements
The collector and the lead author of this dataset, Yihua Zhang, bears full responsibility for any violation of rights that may arise from the collection of the data included in this research.
B Reproducibility Statement and Detailed Experiment Settings
In this section, we provide detailed instructions on the reproduction of our results in Sec. 4, including the settings of training, the implementation details of the tested machine unlearning methods, and the evaluation details in each unlearning scenario. We denote in Sec. 4, 50 out of the 60 styles proposed in UnlearnCanvas are studied and evaluated, with the other 10 styles being intentionally reserved for future explorations, such as generalizations.
B.1 Finetuning Style and Object Classifier with UnlearnCanvas
Style and object classifiers need to be trained as part of the testbed proposed in our evaluation pipeline (Fig. 5). Here, we adopted a ViT-L/16 model [68] pretrained on ImageNet and finetune it on UnlearnCanvas. UnlearnCanvas are split into the train set and test set with a ratio of $9:1$. After hyper-parameter tuning, the classifiers are trained with Adam optimizer at a learning rate of $0.01$ for 10 epochs.
B.2 Finetuning Stable Diffusion with UnlearnCanvas
@@
Practically, we finetune the pretrained Stable Diffusion (SD) v1.5 on UnlearnCanvas for 20k steps with a learning rate of $61e-$ 6. Unless otherwise stated, we strictly follow the training configurations used in Stable Diffusion [1]. For each image in UnlearnCanvas, we annotate the data with text prompt An image of {object} in {style}, where the object and style are the corresponding object and style label. For seed images, the style label we use is ‘photo’ style. We use the training scripts provided by Diffuser official tutorial (https://huggingface.co/docs/diffusers/v0.13.0/en/training/text2image) and the pretraining model card is runwayml/stable-diffusion-v1-5. During training, the checkpoints will be saved every 1000 steps.
@@
To evaluate the quality of the saved checkpoints and selcect the best one for unlearning study, the checkpoints are first used to generate an image set with the same prompt as training (An image of {object} in {style}) by traversing all the possible style and object labels. Each prompt are used to generate $5$ images with different random seed. Each image are sampled with 100 steps with a guidance coeeficient of $9$. The image set for each checkpoint are fed into the style and object classifier trained in Appx. B.1. The model with the highest average performance on all the styles and objects are selected as the testbed for MU study. The classification performance are also used as a reference for later IRA/CRA comparison, which are disclosed in the first row of Fig. 6 (left).
Computing resource.
In this work, we employ 40×40$\times$40 × NVIDIA RTX A6000 GPUs to conduct all the model training, unlearning, image generation, and evaluations. When we finetuned the StableDiffusion on UnlearnCanvas, 8×8$\times$8 × GPUs were used for parallel computing. Other experiments were all carried out in a single-GPU environment. Around 60,000 GPU hours in total were spent to complete all the experiments.
B.3 Implementation of DM Unlearning Methods Studied in This Work
In this work, we inspected a series of stateful MU methods for DMs. For each method, we use their publicly released source codes as code bases, which are listed below:
@@ B.3 Implementation of DM Unlearning Methods Studied in This Work
  • ESD: Based on the suggestions from the authors, ESD is used to only finetune the cross attention-related model weights (ESD-x). Other settings strict follow the ones used in the paper.
@@
  • UCE: This method requires a guided concept (prompt) for each unlearning concept. For style unlearning, we use the prompt An image in {style*} as the guided concept, where style* represents the next style in UnlearnCanvas in alphabetical order. Similarly, An image of object* is used for object unlearning, where object* is the next object in UnlearnCanvas in alphabetical order.
@@
  • SEOT: To generate unlearned images, we use the prompt An {object*} image in {style*}. We then suppress either object* or style* individually. Other settings strict follow the ones used in the paper.
  • SPM: Following the hyperparameters provided by the authors, we trained and obtained Pre-tuned SPMs for all object* and style*. During image generation, we combine the pre-tuned SPMs with the DM. By calculating the association between words in the prompt and the target word, we determine whether to allow the specified word to preserve, and generate the corresponding image.
  • EDiff: Based on the authors’ suggestions, EDiff is used to finetune only the cross-attention-related model weights (EraseDiff-x). During the unlearning process, we adjusted the hyperparameters, specifically the learning rate and the number of unlearning epochs. The model is trained with $5$ epochs.
  • SHS: SHS consists of two stages: trimming and repairing. During the trimming stage, certain weights are re-initialized. The repairing stage then restores the model’s utility. Throughout the unlearning process, we finetuned the hyperparameters, focusing on the learning rate and the number of unlearning epochs. Ultimately, we selected $2$ epochs as the optimal number.
B.4 Evaluation Details of the Adversarial Prompt Generation for Unlearning Robustness
In Sec. 4, we evaluated the robustness of different MU methods against adversarial prompts. Here, we use the state-of-the-art method, UnlearnDiffAtk [59] to generate adversarial prompts. We set the prepended prompt perturbations by $N=5$ tokens for both style and object unlearning. Following the original attack setting in UnlearnDiffAtk [59], to optimize the adversarial perturbations, we sample 50 diffusion time steps and perform PGD running for 40 iterations with a learning rate of 0.01 at each step. Prior to projection onto the discrete text space, we utilize the AdamW optimizer.
B.5 Experiment Details of the Style-Object Combination Unlearning
Unlearning targets.
In Sec. 4, we evaluate the capability of different MU methods on performing unlearning at a finer scale, and we use the style-object combinations as unlearning targets for evaluation. Ideally, the UnlearnCanvas dataset can generate 1200 ($60\times 20$) style-object combinations. In this work, we randomly select 50 of these combinations for evaluation and we list these combinations below. For each method, the same hyper-parameters are used for each MU method as the ones for style and object unlearning in Tab. 2. The unlearning targets include:
  • ‘An image of Architectures in Abstractionism style.’
  • ‘An image of Bears in Artist Sketch style.’
  • ‘An image of Birds in Blossom Season style.’
  • ‘An image of Butterfly in Bricks style.’
  • ‘An image of Cats in Byzantine style.’
  • ‘An image of Dogs in Cartoon style.’
  • ‘An image of Fishes in Cold Warm style.’
  • ‘An image of Flame in Color Fantasy style.’
  • ‘An image of Flowers in Comic Etch style.’
  • ‘An image of Frogs in Crayon style.’
  • ‘An image of Horses in Cubism style.’
  • ‘An image of Human in Dadaism style.’
  • ‘An image of Jellyfish in Dapple style.’
  • ‘An image of Rabbits in Defoliation style.’
  • ‘An image of Sandwiches in Early Autumn style.’
  • ‘An image of Sea in Expressionism style.’
  • ‘An image of Statues in Fauvism style.’
  • ‘An image of Towers in French style.’
  • ‘An image of Trees in Glowing Sunset style.’
  • ‘An image of Waterfalls in Gorgeous Love style.’
  • ‘An image of Architectures in Greenfield style.’
  • ‘An image of Bears in Impressionism style.’
  • ‘An image of Birds in Ink Art style.’
  • ‘An image of Butterfly in Joy style.’
  • ‘An image of Cats in Liquid Dreams style.’
  • ‘An image of Dogs in Magic Cube style.’
  • ‘An image of Fishes in Meta Physics style.’
  • ‘An image of Flame in Meteor Shower style.’
  • ‘An image of Flowers in Monet style.’
  • ‘An image of Frogs in Mosaic style.’
  • ‘An image of Horses in Neon Lines style.’
  • ‘An image of Human in On Fire style.’
  • ‘An image of Jellyfish in Pastel style.’
  • ‘An image of Rabbits in Pencil Drawing style.’
  • ‘An image of Sandwiches in Picasso style.’
  • ‘An image of Sea in Pop Art style.’
  • ‘An image of Statues in Red Blue Ink style.’
  • ‘An image of Towers in Rust style.’
  • ‘An image of Waterfalls in Sketch style.’
  • ‘An image of Architectures in Sponge Dabbed style.’
  • ‘An image of Bears in Structuralism style.’
  • ‘An image of Birds in Superstring style.’
  • ‘An image of Butterfly in Surrealism style.’
  • ‘An image of Cats in Ukiyoe style.’
  • ‘An image of Dogs in Van Gogh style.’
  • ‘An image of Fishes in Vibrant Flow style.’
  • ‘An image of Flame in Warm Love style.’
  • ‘An image of Flowers in Warm Smear style.’
  • ‘An image of Frogs in Watercolor style.’
  • ‘An image of Horses in Winter style.’
Evaluation.
The evaluation of the style-object combination unlearning concerns four quantitative metrics, one for unlearning effectiveness and three for retainability. Before the evaluation, an answer set will be generated exactly following the same procedure introduced in Sec. 3 and Fig. 5 after unlearning each target. First, the UA (unlearning accuracy) will be evaluated for each answer set, which stands for the ratio of images generated by the target prompt that are neither classified into the target object nor the target style class. A high UA denotes a better ability to successfully unlearn the target combination. Second, the retainability of generation associated with those prompts close to the unlearning target prompt will be evaluated. These prompts can be divided into two groups, the ones sharing the same style but not the object class and the ones sharing the object but not the style class. The classification accuracy of the former corresponds to the retainability of the style, i.e., style consistency (SC), while the latter one denotes the object consistency (OC). These two quantitative metrics evaluate how well the unlearning method precisely define the unlearning scope and retain the generation ability of those close but innocent concept. Thirdly, the retainability of the rest unrelated prompts (UP) are evaluated, which is the last quantitative evaluation metric. The results reported in Tab. 3 are averaged over all the unlearning cases shown above.
B.6 Experiment Details of Sequential Unlearning
In Sec. 4, we also evaluated the MU methods with the task of sequential unlearning (SU), where the efficacy of MU methods in handling multiple sequential unlearning requests $\{\mathcal{T}_{i}\}$ are evaluated. This requires models not only to unlearn new targets effectively but also to maintain the unlearning of previous targets, while retaining all other knowledge. In the experiments, 6 styles are randomly selected as the unlearning targets and excluded from the RA evaluation. The UA of all the already unlearned target will be assessed each time a new request is accomplished. The selected 6 styles include:
  • Abstractionism
  • Byzantine
  • Cartoon
  • Cold Warm
  • Ukiyoe
  • Van Gogh
After each unlearning request, the unlearning effectiveness and retainability are evaluated. Specifically, the unlearning accuracy of all the unlearning targets in the previous requests are evaluated to evaluate how the unlearning effect lasts when new unlearning requests arrive. In the meantime, the retainability of all the other concepts that are not selected as unlearning targets are evaluated, and to ease the presentation, the retain accuracy of all the concepts (styles and objects) are averaged and reported.
B.7 Metrics Summary in All Unlearning Settings
Besides the UnlearnCanvas dataset, the various quantitative evaluation metrics proposed in this work are part of the major contributions to a comprehensive and precise evaluation for DM unlearning methods. As there are various unlearning scenarios studied in this work, we provide a summary of these metrics in Tab. A1, including their abbreviations, descriptions, and related tables or figures.
Table A1: A summary of the quantitative metrics used in this work, including their abbreviations, meanings and where they are used.
Metrics Description Usages (Table & Figure)
Style/Object Unlearning
UA Unlearning accuracy Fig. 1, Tab. 1
IRA In-domain unlearning accuracy Fig. 1, Tab. 1
CRA Cross-domain unlearning accuracy Fig. 1, Tab. 1
Unlearning Robustness against Adversarial Prompts
Rob. Unlearning robustness, unlearning accuracy in the presence of adversarial prompts Fig. 1
Style-Object Combination Finer-Scale Unlearning
FU/UA Unlearning accuracy in finer-scale unlearning Fig. 1, Tab. 3
SC Retainability evaluation of style consistency Tab. 3
OC Retainability evaluation of object consistency Tab. 3
UP Retainability evaluation of unrelated prompts Tab. 3
FR Retainability evaluation in finer-scale unlearning, averaged by SC, OC, and UP Fig. 1
Sequential or Continual Unlearning
SU or CU Unlearning accuracy in the context of sequential unlearning Fig. 1, Tab. A5
SR or CR Retainability in the context of sequential unlearning Fig. 1, Tab. A5
C A Detailed Comparison between UnlearnCanvas and WikiArt
Figure A1: Image examples with the same style label from WikiArt [66] and UnlearnCanvas. Images of the same artistic style in UnlearnCanvas exhibit high stylistic consistency compared to WikiArt.
UnlearnCanvas vs. WikiArt.
To the best of our knowledge, WikiArt [73] is the most relevant baseline dataset to ours. In Tab. A2, we provide a direct comparison of the key attributes of these two datasets. UnlearnCanvas differs from WikiArt in the following aspects.
First, UnlearnCanvas includes a greater number of high-resolution images ($15$M) compared to WikiArt ($2$M), a factor that may enhance the training of state-of-the-art DMs.